OpenStudy (anonymous):

how do i convert 0.00000000263 into scientific notation

OpenStudy (anonymous):

so 10^1=10 10^0=1 10^-1=.1 10^-2=.01

OpenStudy (anonymous):

and 2*10^-2=.02

OpenStudy (anonymous):

i think you can figure out the rest based on this?

OpenStudy (anonymous):

its backwards hon

OpenStudy (anonymous):

you have ir right but 263 should be in the front ok

OpenStudy (pagen13):

To write a number in scientific notation: Put the decimal after the first digit and drop the zeroes. To find the exponent count the number of places from the decimal to the end of the number.

jhonyy9 (jhonyy9):

can you rewrite it now so ?

OpenStudy (pagen13):

Actually 2.63 * 10^

jhonyy9 (jhonyy9):

but ten need having exponent

OpenStudy (pagen13):

Yes you would now count how many zeros are in front of 2.63

jhonyy9 (jhonyy9):

yes right so how many will be ?

OpenStudy (pagen13):

Entered problem: Convert 0.00000000263 to scientific notation. The solution is: 0.00000000263=2.63⋅10−9 Explanation In order to write number 0.00000000263 in scientific notation we need to move the decimal point from its current location (black dot) to the new position (red dot). 0.000000002.63 So, we need to move decimal point 9 places to the left. This means that the power of 10 will be negative 9. Now we have that the Number part = 2.63 and Exponent part = -9 So, the solution is: 2.63⋅10−9

jhonyy9 (jhonyy9):

so will be 2,63*10^(-9) yes ?

OpenStudy (pagen13):

Yes